# MMLongBench: Benchmarking Long-Context Vision-Language Models Effectively and Thoroughly

### v1

@@ Abstract
The rapid extension of context windows in large vision–language models has given rise to long-context vision-language models (LCVLMs), which are capable of handling hundreds of images with interleaved text tokens in a single forward pass. In this work, we introduce MMLongBench, the first benchmark covering a diverse set of long-context vision-language tasks, to evaluate LCVLMs effectively and thoroughly. MMLongBench is composed of 13,331 examples spanning five different categories of downstream tasks, such as Visual RAG and Many-Shot ICL. It also provides broad coverage of image types, including various natural and synthetic images. To assess the robustness of the models to different input lengths, all examples are delivered at five standardized input lengths (8K-128K tokens) via a cross-modal tokenization scheme that combines vision patches and text tokens. Through a thorough benchmarking of 46 closed-source and open-source LCVLMs, we provide a comprehensive analysis of the current models’ vision-language long-context ability. Our results show that: i) performance on a single task is a weak proxy for overall long-context capability; ii) both closed-source and open-source models face challenges in long-context vision-language tasks, indicating substantial room for future improvement; iii) models with stronger reasoning ability tend to exhibit better long-context performance. By offering wide task coverage, various image types, and rigorous length control, MMLongBench  [^1] The code and data are available at https://github.com/EdinburghNLP/MMLongBench. provides the missing foundation for diagnosing and advancing the next generation of LCVLMs.
@@ 1 Introduction
Recent advances in long-context modeling unlocked a wide array of new capabilities for both large language models [LLMs; 1, 2] and large vision–language models [3, 4, LVLMs;]. In particular, long-context vision–language models (LCVLMs) represent an important step forward by enabling LVLMs to process hundreds of images and thousands of interleaved text tokens in a single forward pass. This allows applications such as document-level visual question answering [5], multi-hop reasoning across web pages [6], and instruction following grounded in complex visual contexts [7, 8].
@@
  • Insufficient coverage of image types: Most existing benchmarks focus narrowly on either natural images [11, 16] – such as photographs of everyday scenes, objects, or people – or synthetic images [5, 17], such as scanned documents, web pages, or app screenshots. This narrow focus leads to an incomplete understanding of model performance across diverse image types.
  • Lack of context length control: Existing benchmarks miss a consensus on cross-modality length control, especially image tokens. For example, MM-NIAH [18] follows InternVL1.5 [19] to compute text and image tokens together, but other works, such as Visual Haystack [16] and LongDocURL [17], only report the number of images as the context length. This inconsistency makes it difficult to compare model performance across different benchmarks.
  • Single context length: Many long-context benchmarks with text-only inputs [12, 20] standardize context lengths to a few values (e.g., 8K, 32K, 128K) and provide each example with contexts at each standard length. Hence, model developers can easily know the performance change with different lengths. However, such practice is not followed in LCVLM evaluations. For example, in MM-NIAH [18], models are evaluated on long essays (such as web pages) with randomly varying lengths, which complicates systematic analysis of context length effects.
Table 1: Comparison of benchmarks for long-context vision-language models: MM-NIAH [18], Visual Haystack [16], MMNeedle [11], MMLongBench-Doc (MMLB-Doc) [5], M-Longdoc [21], LongDocURL [17], and our MMLongBench. Note that “Summ” refers to summarization, and “DocVQA” stands for document visual question answering (VQA). “$L$” denotes the number of input tokens, and “Mixed” indicates that the dataset includes both natural and synthetic images.
@@
To enable comprehensive evaluation, in this paper, we introduce MMLongBench, a benchmark covering a diverse set of long-context vision-language tasks across five different categories. Specifically, in addition to multimodal Needle-in-a-haystack (NIAH) and long-document VQA (DocVQA) tasks, we also include visual retrieval-augmented generation (VRAG), many-shot in-context learning (ICL), and Summarization in our benchmark. VRAG examples are drawn from knowledge-based VQA datasets [22, 23] using Wikipedia passages to populate long contexts. ICL examples are image classification problems in four domains [24, 25, 26, 27], which require models to perform on-the-fly classification based on hundreds of in-context examples. In the summarization task, models are required to summarize image-based PDF documents [15, 28]. Overall, our benchmark includes diverse downstream tasks and image types to enable a comprehensive evaluation.
In MMLongBench, we use a unified token-count method which counts image tokens based on the number of patches produced by current vision encoders, followed by a $2\times 2$ pixel unshuffle. This approach is consistent with practices adopted in most recent models, such as Qwen2.5-VL [29] and InternVL3 [30], making it well-suited for long-context evaluation. For standard lengths, we equip all examples with five different lengths of context, ranging from 8K to 128K tokens, enabling a thorough analysis of performance changes as the context length increases. Further, we also ensure all datasets are easily extendable to longer contexts.
@@ 3 Our Benchmark — MMLongBench
3.1 Diverse Real Long-Context Applications for LCVLMs
@@
To build a long context, we insert the gold passage(s) (the passage with the answer) among a large set of distracting passages retrieved from Wikipedia. For ViQuAE, we use gold passages from KILT [71] as it is constructed upon TriviaQA [72]. For InfoSeek, we choose the lead section of the named entity’s Wikipedia page as the gold reference and remove all examples for which the answer cannot be found. Then, we split Wikipedia pages into 100-word passages and incrementally add retrieved passages that do not contain the answer or the named entity as distractors until we reach the given input length $L$. For retrieval, we use the named entity instead of the image itself as the query, because text-based retrieval achieves higher recall and provides harder distractors. In ViQuAE, each example contains a single gold passage, and we insert it at six evenly distributed positions, while in Infoseek, the lead sections often contain hundreds of words, resulting in multiple passages. We randomly shuffle them into three permutations. We use the substring exact match (SubEM) as the metric, following previous work [73]. See more details in Section A.1.
@@
Needle-in-a-haystack (NIAH) measures how well an LCVLM can recall a small but important piece of information embedded within a long sequence of mostly unrelated visual and textual inputs. Those tasks have been widely adopted because they are easy to build (can be procedurally generated with a long corpus) and simple to control (can combine any context length and needle position). For this category, we select multiple tasks from Visual Haystack [VH; 16] and MM-NIAH [18]. VH requires models to retrieve images of target objects (the needle) in an image haystack. It is available in two versions: VH-Single and VH-Multi, for finding a single image and multiple images. MM-NIAH contains retrieval (Ret), counting (Count), and reasoning (Reason) tasks in interleaved text and images; each task features both text and image needles.
@@
In VH, we obtain the needle images and the target objects from the original datasets. Then, we accompany these images with multiple negative distractors until the image haystack reaches a given input length $L$. We report accuracy as the metric following the original work [16]. In MM-NIAH, the haystacks are composed of web documents [74] with interleaved text and images. We include all three tasks of retrieval, counting, and reasoning in our benchmark. Similar to VRAG, we split text in each web document into 100-word passages and add passages and images so that the context achieves the input length. We use SubEM and accuracy to evaluate retrieval and reasoning tasks following the original paper [18]. In the counting task, we use the accuracy of summed needle counts for better robustness. Refer to Section A.2 for more details.
@@
Many-shot in-context learning (ICL) tests the model’s capability to adapt to new multimodal tasks on the fly by observing multiple in-context examples, without requiring any parameter updates. Following prior work on long-context LLMs [20, 75, 76], we focus on image classification datasets with large label spaces. Here, we collect four datasets with diverse domains: Stanford Cars [24] for cars, Food101 [25] for food, SUN397 [77] for scenes, and iNat2021 [27] for species. We adjust the number of shots to control the input length $L$, and the number of exemplars in each class is balanced. The 128K context window can accommodate approximately 500 images. To ensure sufficient shots per class, we randomly sample 50 classes from each dataset. We report accuracy on each dataset.
@@
Long-document VQA assesses the model’s aptitude for answering questions that require reasoning over information dispersed across multiple images and text segments within an extended document. We include commonly adopted datasets for evaluating long-document VQA: SlideVQA [81], MMLongBench-Doc [5], and LongDocURL [17]. For documents longer than input length $L$, we truncate the documents evenly from both sides while keeping the answer pages. For shorter documents, we alternately pad the left and right sides with randomly sampled negative documents up to length $L$. However, the padding documents may occasionally contain information related to the question and potentially change the answer. To ensure the validity of questions, we preface each question with the prompt “Based on the Document <Original Doc ID>, answer the following question.” We follow the metrics used in LongDocURL but remove questions with long answers, thereby avoiding LLM-based answer extraction. We list specific details in Section A.5.
@@ 3.3 Standardized Input Length
The input length $L$ is an important factor to consider when we evaluate models’ long context ability, as longer inputs can provide more information but also challenge models to filter out distracting information. As aforementioned in Section 3.1, we can control the input length $L$ for each dataset either by adjusting the number of passages, images, or exemplars, or by truncating the PDF-formatted documents. This allows us to present each example in our benchmark under multiple standardized input lengths and better understand how performance changes as the context length increases. Specifically, our benchmark provides five input lengths $L$: 8K, 16K, 32K, 64K, and 128K tokens, using binary prefixes $K=2^{10}$, and the input length can be easily extended beyond 128K if needed.
@@ 4 Evaluation and Analysis
With broad task coverage, unified token counting, and standardized input length, we are now able to thoroughly examine LCVLMs’ long-context ability across multiple dimensions. In total, we evaluate 46 LCVLMs on MMLongBench. To the best of our knowledge, our evaluation provides the most thorough and controlled comparison of the vision-language long-context ability on broad real-world applications. These models include closed-source models GPT-4o [10], Claude-3.7 [34], and Gemini 2 and 2.5 [84, 43], as well as open-source model families, such as Qwen2.5-VL [29], InternVL3 [30], and Gemma3 [3]. We also consider position extrapolation methods, such as YaRN [37] and V2PE [52] (See Section D.5). We list all models evaluated in Table 10. Following existing works [20], we use greedy decoding for all models for consistency and randomly sample 100 examples from each dataset. More details are in Appendix C.
@@ 4.1 Evaluation on MMLongBench across Tasks and Context Lengths
All models struggle, but closed-source models perform better. Here, we consider the performance at the longest input length of 128K tokens. In general, we observe that all models struggle on our vision-language long-context tasks. For example, even GPT-4o only achieves 62.9 on average, while open-source models perform even worse. We find that Gemini-2.5-Pro stands out as the strongest LCVLMs. Other than ICL, Gemini-2.5-Pro outperforms open-source models by about 20 absolute points. On ICL, although the gap is relatively smaller, due to the strong performance of Qwen2-VL-72B, there is still a difference of about 14 points. While the other closed-source models continue to surpass open-source models, the margin is often under 10 points. Further, Ovis2-34B achieves a score of 41.6 on summarization, similar to GPT-4o (42.4). Qwen2.5-VL-32B achieves a SubEM score of 64.6 on VRAG, even better than Gemini-2.0-Flash. These findings show that while current closed-source models generally perform better, open-source ones are also competitive.
Models can generalize to longer context lengths. Another interesting observation is that some models can generalize to longer context lengths than they are officially designed for. For example, although the context window for Qwen2-VL-72B during training is only 32K tokens, the model can generalize to a 128K input length and still achieve an average score of 51.9. We also observed similar effects on other models, such as Ovis2-34B and InternVL2.5-26B. This phenomenon is likely because the underlying LLMs of those LCVLMs have been trained with longer context windows [29]. We leave further investigation to future work.
Reasoning can improve multimodal long-context ability. We include Gemini-2.0-Flash-T in our evaluation, which is the thinking version of Gemini-2.0-Flash. From the results, we observe that the reasoning ability can consistently improve the Gemini-2.0-Flash on all tasks. While the changes for VRAG, Recall, and ICL are modest, summarization and DocQA exhibit marked improvements of 25.3% and 10.1%, respectively. Then, Gemini-2.5 models exhibit even stronger performance, which are natively designed as thinking models.
Different models exhibit different strengths. Generally, we find that model performance varies considerably across different tasks. For instance, Qwen2.5-VL-32B outperforms InternVL3-38B on VRAG, but underperforms on NIAH. Similarly, Ovis2-34B excels at summarization but struggles on DocQA. These findings further support the necessity of a comprehensive benchmark and low correlation between different task categories. In Sections D.5 and D.6, we also provide additional analysis about the performance of position extrapolations and the lost-in-the-middle phenomenon.
@@
4.2 Can Needle-in-a-Haystack Task Reflect LCVLM’s Overall Long-context Ability?
The needle-in-a-haystack (NIAH) task has been primarily used to evaluate LCVLMs’ long-context abilities. However, it remains unclear whether strong performance on NIAH reliably reflects overall long-context capability on diverse tasks. In this section, we first analyze the difficulty of existing NIAH benchmarks and find that current NIAH tasks are challenging, resulting in limited differentiation between models. Further, we compute Spearman’s rank correlation ($\rho$) between NIAH performance and that on other tasks. Our results show that none of these NIAH tasks consistently correlates with performance across diverse, practical scenarios.
Text-image interleaved NIAH tasks are challenging. In Figure 3, we find that even state-of-the-art models like GPT-4o and Gemini-2.5 struggle to surpass 80% accuracy on VH-Single when the context length is just 8K tokens (approximately 22 images). Most models are just slightly better than a random guess (50%). This demonstrates that locating objects in a large set of images is still hugely challenging for current LCVLMs. See more discussion in Section D.1. Then, we plot the performance of different models on the retrieval, counting, and reasoning tasks of MM-NIAH against their performance on long-document VQA in Figure 2. We find that most models achieve low performance on the counting and reasoning tasks, with scores below 30 and 40, respectively. The difficulty of the tasks and low performance result in poor separability between models. While the retrieval is an easier task, it still does not align well with DocVQA tasks. In short, we find that both VH and MM-NIAH present significant challenges to current LCVLMs, thus showing limited differentiation between models and weak alignment with other tasks.

### v2

@@ Abstract
The rapid extension of context windows in large vision-language models has given rise to long-context vision-language models (LCVLMs), which are capable of handling hundreds of images with interleaved text tokens in a single forward pass. In this work, we introduce MMLongBench, the first benchmark covering a diverse set of long-context vision-language tasks, to evaluate LCVLMs effectively and thoroughly. MMLongBench is composed of 13,331 examples spanning five different categories of downstream tasks, such as Visual RAG and Many-Shot ICL. It also provides broad coverage of image types, including various natural and synthetic images. To assess the robustness of the models to different input lengths, all examples are delivered at five standardized input lengths (8K–128K tokens) via a cross-modal tokenization scheme that combines vision patches and text tokens. Through a thorough benchmarking of 46 closed-source and open-source LCVLMs, we provide a comprehensive analysis of the current models’ vision-language long-context ability. Our results show that: i) performance on a single task is a weak proxy for overall long-context capability; ii) both closed-source and open-source models face challenges in long-context vision-language tasks, indicating substantial room for future improvement; iii) models with stronger reasoning ability tend to exhibit better long-context performance. By offering wide task coverage, various image types, and rigorous length control, MMLongBench  [^1] The code and data are available at https://github.com/EdinburghNLP/MMLongBench. provides the missing foundation for diagnosing and advancing the next generation of LCVLMs.
@@ 1 Introduction
Recent advances in long-context modeling have unlocked a wide array of new capabilities for both large language models [LLMs; 1, 2] and large vision–language models [LVLMs; 3, 4]. In particular, long-context vision–language models (LCVLMs) represent an important step forward by enabling LVLMs to process hundreds of images and thousands of interleaved text tokens in a single forward pass. This allows applications such as document-level visual question answering [5], multi-hop reasoning across web pages [6], and instruction following grounded in complex visual contexts [7, 8].
@@
  • Insufficient coverage of image types: Most existing benchmarks are restricted to either natural images [11, 16] — such as photographs of everyday scenes, objects, or people — or synthetic images [5, 17], such as scanned documents, web pages, or app screenshots. This limited focus leads to an incomplete understanding of model performance across diverse image types.
  • Lack of context length control: Existing benchmarks miss a consensus on cross-modality length control, especially image tokens. For example, while MM-NIAH [18] follows InternVL1.5 [19] to count text and image tokens together, other works, such as Visual Haystack [16] and LongDocURL [17], only report the number of images as the context length. This inconsistency makes it difficult to compare model performance across different benchmarks.
  • Single context length: Many long-context benchmarks with text-only inputs standardize context lengths to a few values (e.g., 8K, 32K, 128K) and provide each example with multiple contexts at those lengths [12, 20]. Hence, model developers can easily know the performance change with different lengths. However, such practice is not followed in LCVLM evaluations. For example, in MM-NIAH [18], models are evaluated on long essays (such as web pages) with randomly varying lengths, which complicates systematic analysis of context length effects.
Table 1: Comparison of benchmarks for LCVLMs: MM-NIAH [18], Visual Haystack [16], MMNeedle [11], MMLongBench-Doc (MMLB-Doc) [5], M-Longdoc [21], LongDocURL [17], and our MMLongBench. “Summ” and “DocVQA” refer to summarization and long-document visual question answering (VQA). “Mixed” indicates both natural and synthetic images. “$L$” denotes the number of input tokens, and “$L$ Control” means counting both text and image tokens together.
@@
To enable comprehensive evaluation, in this paper, we introduce MMLongBench, a benchmark covering a diverse set of long-context vision-language tasks across five different categories. Specifically, in addition to multimodal needle-in-a-haystack (NIAH) and long-document VQA (DocVQA), we also include visual retrieval-augmented generation (VRAG), many-shot in-context learning (ICL), and summarization in our benchmark. VRAG examples are drawn from knowledge-based VQA datasets [22, 23] using Wikipedia passages to populate long contexts. ICL examples are image classification problems in four domains [24, 25, 26, 27], which require models to perform on-the-fly classification based on hundreds of in-context exemplars. In the summarization task, models are required to summarize image-based PDF documents [15, 28]. Overall, our benchmark includes diverse downstream tasks and image types to enable a comprehensive evaluation.
In MMLongBench, we use a unified token-count method which counts image tokens based on the number of patches produced by current vision encoders, followed by a $2\times 2$ pixel unshuffle. This approach is consistent with practices adopted in most recent models, such as Qwen2.5-VL [29] and InternVL3 [30], making it well-suited for long-context evaluation. For a thorough evaluation of input length, we equip all examples with five standardized lengths of context, ranging from 8K to 128K tokens, enabling a thorough analysis of performance changes as the context length increases. Further, we also ensure all datasets are easily extendable to longer contexts.
@@ 3 Our Benchmark — MMLongBench
3.1 Diverse Long-Context Applications for LCVLMs
@@
To build a long context, we insert the gold passage(s) (the passage with the answer) among a large set of distracting passages retrieved from Wikipedia. For ViQuAE, we use gold passages from KILT [72] since it is constructed upon TriviaQA [73]. For InfoSeek, we choose the lead section of the named entity’s Wikipedia page as the gold reference and remove all examples for which the answer cannot be found in the lead section. Then, we split Wikipedia pages into 100-word passages and incrementally add retrieved passages that do not contain the answer or the named entity as distractors until we reach the given input length $L$. For retrieval, we use the named entity instead of the image itself as the query, because text-based retrieval achieves higher recall and provides harder distractors. In ViQuAE, each example contains a single gold passage, and we insert it at six evenly distributed positions, while in Infoseek, the lead sections often contain hundreds of words, resulting in multiple passages. We randomly shuffle them into three permutations. We use the substring exact match (SubEM) as the metric, following previous work [74]. See more details in Section A.1.
@@
Needle-in-a-haystack (NIAH) measures how well an LCVLM can recall a small but important piece of information embedded within a long sequence of mostly unrelated visual and textual inputs. NIAH tasks have been widely adopted because they are easy to build (can be procedurally generated with a long corpus) and simple to control (can combine any context length and needle position). For this category, we select multiple tasks from Visual Haystack [VH; 16] and MM-NIAH [18]. VH requires models to retrieve images of target objects (the needle) in an image haystack. It is available in two versions: VH-Single and VH-Multi, for finding a single image and multiple images. MM-NIAH contains retrieval (Ret), counting (Count), and reasoning (Reason) tasks in interleaved text and images; each task features both text and image needles.
@@
In VH, we obtain the needle images and the target objects from the original dataset. Then, we accompany these needles with multiple negative distractor images until the image haystack reaches a given input length $L$. We report accuracy as the metric following the original work [16]. In MM-NIAH, the haystacks are composed of web documents [75] with interleaved text and images. We include all three tasks of retrieval, counting, and reasoning in our benchmark. Similar to VRAG, we split text in each web document into 100-word passages and add passages and images so that the context achieves the input length. We use SubEM and accuracy to evaluate retrieval and reasoning tasks following the original paper [18]. In the counting task, we use the accuracy of summed needle counts for better robustness. Refer to Section A.2 for more details.
@@
Many-shot in-context learning (ICL) tests the model’s capability to adapt to new multimodal tasks on the fly by observing multiple in-context exemplars, without requiring any parameter updates. Following prior work on long-context LLMs [20, 76, 77], we focus on image classification datasets with large label spaces. Here, we collect four datasets with diverse domains: Stanford Cars [24] for cars, Food101 [25] for food, SUN397 [78] for scenes, and iNat2021 [27] for species. We adjust the number of shots to control the input length $L$, and the number of exemplars in each class is balanced. The 128K context window can accommodate approximately 500 images. To ensure sufficient shots per class, we randomly sample 50 classes from each dataset. We report accuracy on each dataset.
@@
Long-document VQA (DocVQA) assesses the model’s aptitude for answering questions that require reasoning over information dispersed across multiple images and text segments within an extended document. We include commonly adopted datasets for evaluating long-document VQA: SlideVQA [82], MMLongBench-Doc [5], and LongDocURL [17]. For documents longer than input length $L$, we truncate the documents evenly from both sides while keeping the answer pages. For shorter documents, we alternately pad the left and right sides with randomly sampled negative documents up to length $L$. However, the padding documents may occasionally contain information related to the question and potentially change the answer. To ensure the validity of questions, we preface each question with the prompt “Based on the Document <Original Doc ID>, answer the following question.” We follow the metrics used in LongDocURL but remove questions with long answers, thereby avoiding LLM-based answer extraction. We list specific details in Section A.5.
@@ 3.3 Standardized Input Length
The input length $L$ is an important factor to consider when we evaluate models’ long-context ability, as longer inputs can provide more information but also challenge models to filter out distracting information. As aforementioned in Section 3.1, we can control the input length $L$ for each dataset either by adjusting the number of passages, images, or exemplars, or by truncating the PDF-formatted documents. This allows us to present each example in our benchmark under multiple standardized input lengths and better understand how performance changes as the context length increases. Specifically, our benchmark provides five input lengths $L$: 8K, 16K, 32K, 64K, and 128K tokens, using binary prefixes $K=2^{10}$, and the input length can be easily extended beyond 128K if needed.
@@ 4 Evaluation and Analysis
With broad task coverage, unified token counting, and standardized input length, we are now able to thoroughly examine LCVLMs’ long-context ability across multiple dimensions. In total, we evaluate 46 LCVLMs on MMLongBench. To the best of our knowledge, our evaluation provides the most thorough and controlled comparison of the vision-language long-context ability on broad real-world applications. These models include closed-source models GPT-4o [10], Claude-3.7 [34], and Gemini 2 and 2.5 [85, 43], as well as open-source model families, such as Qwen2.5-VL [29], InternVL3 [30], and Gemma3 [3]. We also consider position extrapolation methods, such as YaRN [37] and V2PE [52] (See Section D.5). The full list of evaluated models is provided in Table 10. Following existing works [20], we use greedy decoding for all models for consistency and randomly sample 100 examples from each dataset. More details on the experimental setup are in Appendix C.
@@ 4.1 Evaluation on MMLongBench across Tasks and Context Lengths
All models struggle, but closed-source models perform better. Here, we consider the performance at the longest input length of 128K tokens. In general, we observe that all models struggle on our vision-language long-context tasks. For example, even GPT-4o only achieves 62.9 on average, while open-source models perform even worse. We find that Gemini-2.5-Pro stands out as the strongest LCVLM. Other than ICL, Gemini-2.5-Pro outperforms open-source models by about 20 absolute points. On ICL, although the gap is relatively smaller, due to the strong performance of Qwen2-VL-72B, there is still a difference of about 14 points. While the other closed-source models continue to surpass open-source models, the margin is often under 10 points. Further, Ovis2-34B achieves a score of 41.6 on summarization, similar to GPT-4o (42.4). Qwen2.5-VL-32B achieves a SubEM score of 64.6 on VRAG, even better than Gemini-2.0-Flash. These findings show that while current closed-source models generally perform better, open-source ones are also competitive.
Models can generalize to longer context lengths. Another interesting observation is that some models can generalize to longer context lengths than they are officially designed for. For example, although the context window for Qwen2-VL-72B during training is only 32K tokens, the model can generalize to a 128K input length and still achieve an average score of 51.9. We also observe similar effects on other models, such as Ovis2-34B and InternVL2.5-26B. This phenomenon is likely because the underlying LLMs of those LCVLMs have been trained with longer context windows [29]. We leave further investigation to future work.
Reasoning can improve multimodal long-context ability. We include Gemini-2.0-Flash-T in our evaluation, which is the thinking variant of Gemini-2.0-Flash. From the results, we observe that the reasoning ability can consistently improve the Gemini-2.0-Flash on all tasks. While the changes for VRAG, Recall, and ICL are modest, summarization and DocVQA exhibit marked improvements of 25.3% and 10.1%, respectively. Then, Gemini-2.5 models exhibit even stronger performance, which are natively designed as thinking models.
Different models exhibit different strengths. Generally, we find that model performance varies considerably across different tasks. For instance, Qwen2.5-VL-32B outperforms InternVL3-38B on VRAG, but underperforms on NIAH. Similarly, Ovis2-34B excels at summarization but struggles on DocVQA. These findings further support the necessity of a comprehensive benchmark covering diverse downstream tasks. In Sections D.5 and D.6, we also provide additional analysis about the performance of position extrapolations and the lost-in-the-middle phenomenon.
@@
4.2 Can Needle-in-a-Haystack Tasks Reflect LCVLM’s Overall Long-Context Ability?
The needle-in-a-haystack (NIAH) has been primarily used to evaluate LCVLMs’ long-context abilities. However, it remains unclear whether strong performance on NIAH reliably reflects overall long-context capability on diverse tasks. In this section, we first analyze the difficulty of existing NIAH benchmarks and find that current NIAH tasks are challenging, resulting in limited differentiation between models. Further, we compute Spearman’s rank correlation ($\rho$) between NIAH performance and that on other tasks. Our results show that none of these NIAH tasks consistently correlates with performance across diverse, practical scenarios.
Text-image interleaved NIAH tasks are challenging. In Figure 3, we find that even state-of-the-art models like GPT-4o and Gemini-2.5 struggle to surpass 80% accuracy on VH-Multi when the context length is just 8K tokens (approximately 22 images). Most models are just slightly better than a random guess (50%). This demonstrates that locating objects in a large set of images is still hugely challenging for current LCVLMs. See more discussion in Section D.1. Then, we plot the performance of different models on the retrieval, counting, and reasoning tasks of MM-NIAH against their performance on long-document VQA in Figure 2. We find that most models achieve low performance on the counting and reasoning tasks, with scores below 30 and 40, respectively. The difficulty of the tasks and low performance result in poor separability between models. While the retrieval is an easier task, it still does not align well with DocVQA tasks. In short, we find that both VH and MM-NIAH present significant challenges to current LCVLMs, thus showing limited differentiation between models and weak alignment with other tasks.